# The ML Supply Chain in the Era of Software 2.0: Lessons Learned from Hugging Face

### v1

@@ 8. Threats to Validity
Construct Validity. We rely on information that is available through the Hugging Face API in a machine readable format. We addressed cases where information was missing from this source by applying Python’s request and BeautifulSoup libraries to retrieve missing information. However, this means that we likely missed licensing and dependency information that was presented in natural language. We measure and mitigate this risk with a review of the top-100 “unlicensed” models sorted by number of downloads, which showed that the majority did not, in fact, declare a license or declared a license in a non-standard or transitive way.
@@
External Validity. The generalizability of our findings is restricted by our choice of Hugging Face, which might not be fully generalizable to all model-sharing platforms. It is possible that we missed other challenges and obstacles faced by other model hubs or that different model-sharing sites have different levels of documentation and dependency management techniques. Additionally, given that the field of AI is rapidly evolving and new models are created every day, the conclusions we glean may be different if we were to take an updated snapshot of the supply chain at some future point. The extent to which this factor might change our results is unknown, but we maintain that it is important to capture an understanding of this nascent supply chain so that future investigations can understand the ways in which the supply chain has changed over time.
@@ 9. Conclusion
In this work we conduct a comprehensive mining study and analysis of the ML supply chain on Hugging Face. We construct a reusable dataset of over 750,000 models and 175,000 datasets, which we make freely available in our replication package (Stalnaker et al., 2025). Using this data we create a supply chain graph that models the relationships between ML models and the models they depend on, describe the notable features and properties of this supply chain, analyze the current state of documentation on Hugging Face, and consider the current state of licensing and license compliance in the ecosystem. We offer a rich discussion of the challenges that not just Hugging Face but any sizable model hub will encounter. Lastly, we offer concrete and actionable recommendations that can be adopted by any model hosting platform to facilitate better supply chain management.
@@
We provide an anonymized replication package containing the raw mined data, our cleaned and normalized dataset, scripts for data cleaning, mining utilities, analysis scripts, and other data required for verifiability (Stalnaker et al., 2025).

### v2

@@ 11. Threats to Validity
Construct Validity. We rely on information that is available through the Hugging Face API in a machine-readable format. We addressed cases where information was missing from this source by applying Python’s request and BeautifulSoup libraries to retrieve missing information. However, this means that we likely missed licensing and dependency information that was presented in natural language. We measure and mitigate this risk with a review of the top 100 “unlicensed” models sorted by number of downloads, which showed that the majority did not, in fact, declare a license or declared a license in a non-standard or transitive way.
@@
External Validity. The generalizability of our findings is restricted by our choice of Hugging Face, which might not be fully generalizable to all model-sharing platforms. It is possible that we missed other challenges and obstacles faced by other model hubs or that different model-sharing sites have different levels of documentation and dependency management techniques, given our assumption that highly-downloaded models tend towards better documentation. We also observed that relatively few models and datasets declared licensing information, and models often did not declare any datasets. This implies that a great portion of the ecosystem may yet be undocumented. As such, our results may be biased towards the components that do provide documentation, and may be less generalizable to those which do not. We attempt to mitigate this risk by providing a detailed analysis of top-rated components to capture the specific challenges facing components that are used frequently and are thus applicable to many users. Additionally, given that the field of AI is rapidly evolving and new models are created every day, the conclusions we reached might be different if we were to take an updated snapshot of the supply chain at some future point. The extent to which this factor might change our results is unknown, but we maintain that it is important to capture an understanding of this nascent supply chain so that future investigations can understand the ways in which the supply chain has changed over time.
@@ 12. Conclusion
In this work we conduct a comprehensive mining study and analysis of the ML supply chain on Hugging Face. We construct a reusable dataset of over 750,000 models and 175,000 datasets, which we make freely available in our replication package (Stalnaker et al., 2025). Using this data we create a supply chain graph that illustrates the relationships between ML models and the models they depend on, describe the notable features and properties of this supply chain, analyze the current state of documentation on Hugging Face, and consider the current state of licensing and license compliance in the ecosystem. We offer a rich discussion of the challenges that not just Hugging Face but any sizable model hub will encounter. Lastly, we offer concrete and actionable recommendations for various stakeholders involved with the ML supply chain that can be adopted to facilitate better supply chain management.
@@
We provide a replication package containing the raw mined data, our cleaned and normalized dataset, scripts for data cleaning, mining utilities, analysis scripts, and other data required for verifiability (Stalnaker et al., 2025).